# MACTAS: Self-Attention-Based Module for Inter-Agent Communication in Multi-Agent Reinforcement Learning

### v2

@@ Abstract.
Communication is essential for the collective execution of complex tasks by human agents, motivating interest in communication mechanisms for multi-agent reinforcement learning (MARL). However, existing communication protocols in MARL are often complex and non-differentiable. In this work, we introduce a self-attention-based communication module that exchanges information between the agents in MARL. Our proposed approach is fully differentiable, allowing agents to learn to generate messages in a reward-driven manner. The module can be seamlessly integrated with any action-value function decomposition method and can be viewed as an extension of such decompositions. Notably, it includes a fixed number of trainable parameters, independent of the number of agents. Experimental results on the SMAC and SMACv2 benchmarks demonstrate the effectiveness of our approach, which achieves state-of-the-art performance on a number of maps.
Key words and phrases:
Multi-Agent Reinforcement Learning, Cooperation in Multi-Agent Reinforcement Learning, Communication in Multi-Agent Reinforcement Learning
1. Introduction
Multi-agent reinforcement learning (MARL) Ning and Xie (2024); Du and Ding (2021); Nguyen et al. (2020); Oroojlooy and Hajinezhad (2023); Wong et al. (2023) promises efficient control in large-scale systems through the collective learning of many agents in dynamic environments. Multiple agents interact within a shared environment to learn optimal policies. MARL is defined as the process in which agents optimize their behaviors based on the history of their interactions with the environment, i.e., on their experience. Research in this area aims to design methods that enable agents to learn quickly, i.e., from minimal experience. The challenges and complexities in using MARL for controlling multi-agent systems include coordination among agents, non-stationary environments due to evolving agent policies, and scalability issues. Advanced MARL approaches often involve sophisticated techniques, including decentralized training, communication protocols among agents, and hierarchical methods to manage multi-agent interactions as effectively as possible. Theoretical advancements and practical applications of MARL demonstrate its potential in diverse fields ranging from robotics Orr and Dutta (2023) and autonomous driving Yadav et al. (2023); Schmidt et al. (2022); Althamary et al. (2019) to telecommunications Li et al. (2022b); Feriani and Hossain (2021).
Figure 1. 3s5z_vs_3s6z scenario in the StarCraft Multi-Agent Challenge (SMAC).
The most prominent research topics in MARL are the action-value function decompositions Wang et al. (2021); Liu et al. (2023b), where agents share a joint reward according to their contributions, and inter-agent communication Liu et al. (2023a); Guo et al. (2023), where agents exchange information in order to coordinate their joint policy. The most challenging multi-agent environments are partially observable, typically by limiting agents’ information to their close proximity, and the achievement of their common goals is predicated upon coordination, which needs to involve some form of communication. In most research, the action-value function and inter-agent communication are issues addressed separately. Therefore, in this paper, we introduce communication into the very core of the action-value function decomposition.
The prevailing paradigm in cooperative MARL is Centralized Training and Decentralized Execution (CTDE), where the training algorithm utilizes the global state, but agents can only use their own local observations. We argue that in most challenging MARL tasks, this is insufficient. Moreover, in most challenging MARL tasks, some form of inter-agent communication is a necessity, which can be provided relatively easily (e.g., through WiFi). Therefore, we introduce a communication protocol that collects information from each agent in a central entity and broadcasts the combined information to all agents. Because information aggregation is based on attention, our proposed architecture requires a constant number of weights, plus an embedding, for each agent. It is easy to apply using readily available and efficient implementations of the Transformer encoder layer.
The contribution of this paper may be summarized as follows:
  • We introduce Multi-Agent Communication via Transformer on Agents’ States (MACTAS) module for inter-agent communication in MARL. This module allows the agent to share their state of consciousness.
  • We verify MACTAS experimentally on several challenging maps from SMACSamvelyan et al. (2019) and SMACv2 Ellis et al. (2023), giving excellent results and establishing state-of-the-art in some cases.
2. Related Work
Single-Agent Reinforcement Learning (SARL) Sutton and Barto (2018) addresses trial-and-error learning in sequential decision-making under uncertainty. In Multi-Agent Reinforcement Learning (MARL) Ning and Xie (2024), a group of agents learns to act in the same dynamic environment, collectively modifying its state. There are various settings for MARL, depending on whether the agents operate synchronously or asynchronously, whether they have common or adverse goals, and other circumstances. In this paper, we focus on the synchronous and cooperative MARL, with a single reward value for all agents at every time instance.
MARL has been addressed by extending well-established SARL algorithms, including Multi-Agent SAC Wei et al. (2018), Multi-Agent PPO Yu et al. (2022), and Multi-Agent DDPG Lowe et al. (2017).
Methods of cooperative MARL often adopt a centralized training with decentralized execution (CTDE) paradigm, in which agents share information during training but act based solely on their own observations during execution. However, this paradigm suffers from the global action-value function, whose complexity grows exponentially with the number of agents Kraemer and Banerjee (2016). Therefore, action-value function decomposition is a fundamental problem in MARL. Existing decomposed architectures include modules assigned to agents. These modules are fed with agents’ observations and their previous actions and produce agents’ contributions to the architecture’s output. The best action for the agent is the one that makes the highest contribution. In VDN Sunehag et al. (2018), these contributions are just added. In QMIX Rashid et al. (2018), a neural module, the mixing network, produces its weighted sum with positive weights. In Weighted QMIX Rashid et al. (2020), the architecture is trained with a special weight assigned to the collective actions deemed the best. In QTRAN Son et al. (2019), an additional module fed with all the actions of all agents corrects the weighted sum of the agents’ contributions. In Qatten Yang et al. (2020), the mixing network is based on attention. In CDS Li et al. (2021), a component of the mixing network optimizes the amount of exploration undertaken by the agents. In DVD Li et al. (2022a), the internal state of the modules assigned to agents is also passed to the mixing networks. In SHAQ Wang et al. (2022), the mixing network is enriched by several components based on Shapley value theory. In QPLEX Wang et al. (2021), a carefully engineered duplex dueling of pairs of contributions of agents is introduced into the mixing network. In NA2Q Liu et al. (2023b), a similar duplex dueling is performed with the use of attention.
@@ 2. Related Work
Another fundamental problem in cooperative MARL is the coordination of agents’ activity for which a certain communication protocol is required. A number of solutions have been proposed for this purpose: CommNet Sukhbaatar et al. (2016) introduces differentiable communication by aggregating a learned broadcast vector from each agent, TarMAC Das et al. (2019) uses signature-based soft attention to direct agent communication towards recipients, NDQ Wang et al. (2019) applies two forms of regularization – mutual information maximization and message entropy minimization to optimize message content, GA-Comm Liu et al. (2020) uses two-stage attention mechanism to limit number of messages exchanged between agents, IS Kim et al. (2021) encodes trajectory of future actions for each agent and shares them with others, in MAIC Yuan et al. (2022) agents explicitly model their teammates and use it to bias their value function, DHCG Liu et al. (2023a) learns a dependency graph between agents and uses it to limit the number of communication edges, TEM Guo et al. (2023) uses a Transformer to simulate communication resembling email exchange, where agents only directly communicate with teammates nearby, but can forward messages from one to another, ACE Li et al. (2023) proposes bidirectional action-dependent Q-learning.
@@
We acknowledge various applications of attention or Transformer architectures in MARL algorithms, such as in Qatten mixing network, or in communication algorithms like NA2Q, GA-Comm, and TEM. None of the previous works has applied the Transformer architecture directly as a communication protocol. Our work fills this gap. We consider communication methods that can be coupled with an arbitrary action-value decomposition method. We also allow agents to send messages based only on their local information (current observation, historical observations, and historical actions). In this framework, the best results so far have been achieved by MAIC Yuan et al. (2022).
3. Method
Figure 2. The proposed MACTAS architecture.
3.1. Problem definition
@@
In this paper, we discuss the problem of MARL using the formalism of a cooperative Decentralized Partially Observable Markov Decision Process (Dec-POMDP) with communication. Dec-POMDP is defined by a tuple $\langle\mathbb{N},\mathbb{S},\mathbb{A},\mathcal{R},\mathcal{P},\Omega,\mathcal{O},\gamma\rangle$, where $\mathbb{N}=\{1,\dots,n\}$ represents the team of agents, $\mathbb{S}$ is the global environmental state space, $\mathbb{A}=\mathbb{A}^{1}\times\dots\times\mathbb{A}^{n}$ is the joint action space, $\mathcal{R}$ is the reward function, $\mathcal{P}$ is the state transition probability function, $\Omega=\Omega^{1}\times\dots\times\Omega^{n}$ is the joint observation space, $\mathcal{O}$ is the observation function and $\gamma$ is the discount factor. At time $t=1,2,\dots$ each $i$-th agent makes an observation, $o^{i}_{t}=\mathcal{O}^{i}(s_{t}),o^{i}_{t}\in\Omega^{i}$, of the environment state $s_{t}\in\mathbb{S}$. Agents may exchange messages with others through a certain channel that may generally refine these messages with some processing.
@@
Each agent maintains a representation, $\mathcal{I}^{i}_{t}$, of observations it has made, messages it has received, and actions it has taken. Based on thereof and a policy, $\pi^{i}$, the agent performs an action, $a^{i}_{t}\in\mathbb{A}^{i}$, $a^{i}_{t}\sim\pi^{i}(\cdot|\mathcal{I}^{i}_{t})$. The team action $a_{t}=[(a^{1}_{t}),\dots,(a^{n}_{t})]$ impacts the next environment state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$ and the reward $r_{t}=\mathcal{R}(s_{t},a_{t},s_{t+1})$ is given collectively to all agents.
@@
The goal of the agents’ learning is to optimize their team policy $\pi=[\pi^{1},\dots,\pi^{n}]$ to expect in each environment state $s$ the highest sum of discounted rewards
$\mathcal{V}^{\pi}(s)=\mathbb{E}\left(\sum_{i\geq 0}\gamma^{i}r_{t+i}\Big|s_{t}=s,\text{policy in use}=\pi\right),$
where $\gamma\in[0,1]$.
@@
The goal of this paper is to establish the message exchange protocol that supports the above goal of learning.
3.2. MACTAS
@@
Intuitively, we understand the $f^{i}$ module as the consciousness of the $i$-th agent. Based on what the agent is aware of, it determines its contribution to future discounted rewards, depending on the action it currently takes. In order to give a global view on the current state to the agent’s consciousness, we introduce a communication mechanism. We divide the $f^{i}$ module after its recurrent part and feed its state $h^{i}_{t}\in\real^{n_{h}}$ to the module $g$, where the states of consciousness of the agents are confronted. As a result, the $g$ module produces increments, $z^{i}_{t}\in\real^{n_{h}}$. The sum $z^{i}_{t}+h^{i}_{t}$ is then passed to the second MLP part of the $f^{i}$ module.
Consequently, the consciousness of each agent is based not only on its observations and previous actions, but also on observations and previous actions of all other agents.
@@
  1. (1)
    It is fed with $n$ vectors and produces $n$ vectors, both of size $\dim(h^{i}_{t})$. $i$-th input vector corresponds to the $i$-th output vector.
  2. (2)
@@
  3. (3)
    We use a separate Adam optimizer Kingma and Ba (2017) for the Transformer parameters; the rest of the architecture is trained with its originally proposed optimizer (typically it is RMSProp Tieleman and Hinton (2012)).
  4. (4)
    The weights of the output layer of the top Transformer module are initialized with zeros. Therefore, initially, this module does not impact the state of consciousness of the agents.
A good performance of the Transformer in this role is an expected result. This architecture has been specifically optimized to transform a set of embeddings into another set of the same size, which is exactly what we require in the $g$ module.
3.3. Implementation considerations
@@
A distributed implementation of a multi-agent system with our proposed module is also possible. Then, each agent needs to take on the part of processing that consists of refining $h^{i}_{t}$ and results in $z^{i}_{t}$. That requires exchange of messages with other agents as many times as many Transformer layers are stacked in MACTAS, each time with $O(n)$ messages; $O(n^{2})$ for all agents. In training, the agents need to aggregate increments in their shared weights.

### v3

# MACTAS: Self-Attention-Based Inter-Agent Communication in Multi-Agent Reinforcement Learning with Action-Value Function Decomposition

@@ Abstract
Communication is essential for the collective execution of complex tasks by human agents, motivating interest in communication mechanisms for multi-agent reinforcement learning (MARL). However, existing communication protocols in MARL are often complex and non-differentiable. In this work, we introduce a self-attention-based communication method that exchanges information between the agents in MARL. Our proposed approach is fully differentiable, allowing agents to learn to generate messages in a reward-driven manner. The method can be seamlessly integrated with any action-value function decomposition algorithm and can be viewed as an orthogonal extension of such decompositions. Notably, it includes a fixed number of trainable parameters, independent of the number of agents, which makes it scalable to large systems. Experimental results on the SMACv2 benchmark demonstrate the effectiveness of our approach, which achieves state-of-the-art performance on a number of maps.
**footnotetext: Equal contribution.
1 Introduction
Multi-agent reinforcement learning (MARL) Ning and Xie (2024); Du and Ding (2021); Nguyen et al. (2020); Oroojlooy and Hajinezhad (2023); Wong et al. (2023) offers a powerful framework for controlling large-scale systems by enabling multiple agents to learn collectively while interacting within a shared, dynamic environment. In MARL, agents seek to optimize their policies based on accumulated experience, derived from the history of their interactions with both the environment and other agents. A central objective of MARL research is to develop algorithms that achieve effective learning with minimal experience, thereby improving sample efficiency.
Despite its promise, MARL presents several fundamental challenges, including the need for coordination among agents, non-stationarity induced by concurrently evolving policies, and scalability to systems with many agents. To address these challenges, advanced MARL methods incorporate a range of techniques such as decentralized or partially centralized training, explicit communication mechanisms, and hierarchical structures that facilitate efficient multi-agent coordination. Ongoing theoretical developments and practical successes highlight the versatility of MARL across a broad spectrum of applications, including robotics Orr and Dutta (2023), autonomous driving Yadav et al. (2023); Schmidt et al. (2022); Althamary et al. (2019), and telecommunications Li et al. (2022b); Feriani and Hossain (2021).
Figure 1: The strategic sacrifice tactics found by MACTAS for the 3s5z_vs_3s6z map from the SMAC benchmark, on which MACTAS outperforms the baselines significantly. The agent at the top of the figure (“Local Player”) lures the opponents away and sacrifices its own life to enable the remaining agents to gain a strategic advantage and win the episode.
Two of the most actively studied directions in MARL are action–value function decomposition methods Wang et al. (2021); Liu et al. (2023b), which allocate a shared team reward according to individual agent contributions, and inter-agent communication mechanisms Liu et al. (2023a); Guo et al. (2023), which enable coordination through information exchange. Many of the most challenging multi-agent environments are only partially observable, often restricting agents’ perceptions to their local neighborhoods. In such settings, successful task completion critically depends on effective coordination, which in turn typically requires some form of communication. In our proposed architecture, communication is integrated directly into the action–value function decomposition itself.
The dominant framework for cooperative MARL is Centralized Training and Decentralized Execution (CTDE), wherein the learning algorithm has access to the global state during training, while each agent relies solely on its local observations at execution time. We contend that this paradigm is inadequate for many complex MARL problems. In practice, reliable inter-agent communication is often feasible—for example, via wireless networks—and can significantly enhance coordination. Motivated by this observation, we introduce a communication mechanism based on self-attention: The latent state of each agent is updated based on other agents’ states, similar to how word embeddings are updated based on other words in the Transformer encoder Vaswani et al. (2017). By leveraging self-attention, the proposed architecture requires only a constant number of shared parameters, along with agent-specific embeddings. Furthermore, it can be efficiently implemented using standard Transformer encoder layers, making it both scalable and practical.
Although self-attention has been employed in different contexts in MARL before, including communication, this is the first successful approach to using self-attention on its own as a communication tool in MARL, without the need for graph-structure learning or global state reconstruction. Moreover, our method can be combined with an arbitrary action-value decomposition. None of the previous works has applied self-attention directly as a communication protocol. Our work fills this gap. It is the first algorithm showing that self-attention on agent internal states is sufficient for effective and reliable communication without the need for explicit teammate modeling or message preparation mechanisms.
The contribution of this paper includes:
  • We introduce Multi-Agent Communication via Transformer on Agents’ States (MACTAS) for inter-agent communication in MARL based on a self-attention mechanism applied directly to internal state embeddings of agents. MACTAS is the first self-attention-based communication mechanism with the number of trainable parameters independent of the number of agents, which makes it scalable to large systems. Moreover, it can be combined with any action-value function mixing architecture.
  • We verify experimentally MACTAS in combination with the most popular decomposition methods on challenging maps from SMACv2, showing results better than current state-of-the-art algorithms in most cases. We also perform experiments on SMAC, Level-based Foraging, and Hallway, proving that MACTAS is more reliable in a diverse set of scenarios.
Our approach outperforms state-of-the-art communication algorithms (including a method that uses self-attention and global state reconstruction) in combination with common mixers. Moreover, our method proves robust on a diverse set of tasks, while each baseline fails on some relatively simple task (like Hallway or the 2c_vs_64zg SMAC map). Our algorithm can learn sophisticated tactics, such as sacrificing one agent to win the entire battle (as shown in Figure 1). A detailed study of the tactics found by our method is presented in Appendix E.
2 Related Work
Single-agent reinforcement learning (SARL) Sutton and Barto (2018) studies trial-and-error learning for sequential decision-making under uncertainty. Multi-agent reinforcement learning (MARL) Ning and Xie (2024) generalizes this framework to settings in which multiple agents simultaneously interact with and influence a shared, dynamic environment. MARL encompasses a wide range of problem formulations, varying along dimensions such as synchronous versus asynchronous decision-making, cooperative versus competitive objectives, and differing information structures. In this work, we restrict our attention to the synchronous, fully cooperative setting, where all agents share a common reward signal at each time step.
A major limitation of the CTDE framework is the exponential growth of the joint action–value function with the number of agents Kraemer and Banerjee (2016). As a result, action–value function decomposition has emerged as a central challenge in MARL. Decomposition-based architectures typically assign each agent a local module that processes its observations and previous actions to produce an individual contribution to the global value. An agent’s optimal action is then determined by maximizing its contribution to the joint value estimate.
Several decomposition approaches have been proposed in the literature. Value Decomposition Networks (VDN) Sunehag et al. (2018) aggregate agent contributions through simple summation, while QMIX Rashid et al. (2018) employs a mixing network that computes a weighted sum with non-negative weights to preserve monotonicity. Weighted QMIX Rashid et al. (2020) further refines this approach by assigning special weights to optimal joint actions during training. QTRAN Son et al. (2019) introduces an auxiliary module that incorporates the joint action to correct the decomposed value. Attention mechanisms are leveraged in Qatten Yang et al. (2020) to construct a more flexible mixing network. Additional extensions include CDS Li et al. (2021), which adjusts agent exploration through the mixing network; DVD Li et al. (2022a), which incorporates agents’ internal states into value aggregation; SHAQ Wang et al. (2022), which augments the mixing network using principles from Shapley value theory; QPLEX Wang et al. (2021), which introduces a duplex dueling architecture over agent contributions; and NA2Q Liu et al. (2023b), which adopts a similar duplex dueling strategy enhanced with attention mechanisms.
@@ 2 Related Work
Another fundamental problem in cooperative MARL is the coordination of agents’ activities, which requires a communication protocol. A number of solutions have been proposed for this purpose: CommNet Sukhbaatar et al. (2016) introduces differentiable communication by aggregating a learned broadcast vector from each agent, TarMAC Das et al. (2019) uses signature-based soft attention to direct agent communication towards recipients, NDQ Wang et al. (2019) applies two forms of regularization – mutual information maximization and message entropy minimization to optimize message content, GA-Comm Liu et al. (2020) uses two-stage attention mechanism to limit number of messages exchanged between agents, IS Kim et al. (2021) encodes trajectory of future actions for each agent and shares them with others, DHCG Liu et al. (2023a) learns a dependency graph between agents and uses it to limit the number of communication edges, TEM Guo et al. (2023) uses a Transformer to simulate communication resembling email exchange, where agents only directly communicate with teammates nearby, but can forward messages from one to another, ExpoComm Li et al. (2025) utilizes exponential graphs to design the communication topology. Also, in MAIC Yuan et al. (2022) agents explicitly model their teammates and use it to bias their value function, MASIA Guan et al. (2022) uses attention-based communication architecture with a decoder module to reconstruct the global state, CommFormer Hu et al. (2024) uses the attention encoder and decoder architecture combined with learning a communication graph, having the number of parameters independent of the number of agents.
@@
Attention and Transformer architectures have been applied in MARL algorithms, such as in Qatten and NA2Q mixing networks, or in communication mechanisms like GA-Comm, TEM, CommFormer, and MASIA. However, our approach is conceptually different: We apply the attention mechanism directly to generate the agents’ messages, without additional graph-structure learning (as in CommFormer) or decoding the global state (as in MASIA). Our proposed use of attention makes the number of trainable parameters independent of the number of agents (which enables deployment at scale), unlike in MASIA. Moreover, our method can be combined with any mixing architecture, unlike CommFormer. We consider communication methods that can be coupled with an arbitrary action-value decomposition method. We also allow agents to send messages based only on their local information (current observation, historical observations, and historical actions). In this framework, MAIC has achieved the best results so far. MASIA is a communication protocol that uses self-attention. Therefore, in the experimental study, we compare our proposed MACTAS to both MAIC and MASIA.
3 Method
In this section, we formalize the problem statement, present the details of applying self-attention in MACTAS, and discuss its distributed deployment.
Figure 2: The proposed MACTAS architecture.
Problem definition.
@@ 3 Method
In this paper, we discuss the problem of MARL using the formalism of a cooperative Decentralized Partially Observable Markov Decision Process (Dec-POMDP) with communication. Dec-POMDP is defined by a tuple $\langle\mathbb{N},\mathbb{S},\mathbb{A},\mathcal{R},\mathcal{P},\Omega,\mathcal{O},\gamma\rangle$, where $\mathbb{N}=\{1,\dots,n\}$ represents the team of agents, $\mathbb{S}$ is the global environmental state space, $\mathbb{A}=\mathbb{A}^{1}\times\dots\times\mathbb{A}^{n}$ is the joint action space, $\mathcal{R}:\mathbb{S}\times\mathbb{A}\times\mathbb{S}\rightarrow\mathbb{R}$ is the reward function, $\mathcal{P}:\mathbb{S}\times\mathbb{A}\rightarrow P(\mathbb{S})$ is the state transition probability function, $\Omega=\Omega^{1}\times\dots\times\Omega^{n}$ is the joint observation space, $\mathcal{O}:\mathbb{N}\times\mathbb{S}\rightarrow\Omega$ is the observation function and $\gamma\in[0,1]$ is the discount factor. At time $t=1,2,\dots$ the $i$-th agent makes an observation, $o^{i}_{t}=\mathcal{O}(i,s_{t}),o^{i}_{t}\in\Omega^{i}$, of the environment state $s_{t}\in\mathbb{S}$. Agents may exchange messages through a channel that generally refines them with some processing.
@@
We define a set of representations $\mathcal{J}=\mathcal{J}^{1}\times\dots\times\mathcal{J}^{n}$. Each agent at timestep $t$ maintains a representation, $\mathcal{I}^{i}_{t}\in\mathcal{J}^{i}$, of observations it has made, messages it has received, and actions it has taken. Based on thereof and a policy, $\pi^{i}:\mathcal{J}^{i}\rightarrow\mathbb{A}^{i}$, the agent performs an action, $a^{i}_{t}\in\mathbb{A}^{i}$, $a^{i}_{t}\sim\pi^{i}(\cdot|\mathcal{I}^{i}_{t})$. The team action $a_{t}=[(a^{1}_{t}),\dots,(a^{n}_{t})]$ impacts the next environment state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$ and the reward $r_{t}=\mathcal{R}(s_{t},a_{t},s_{t+1})$ is given collectively to all agents.
@@
The goal of the agents’ learning is to optimize their team policy $\pi=[\pi^{1},\dots,\pi^{n}]$ to expect in each environment state $s$ the highest sum of discounted rewards:
$\mathcal{V}^{\pi}(s)=\mathbb{E}\left(\sum_{i\geq 0}\gamma^{i}r_{t+i}\Big|s_{t}=s,\text{policy in use}=\pi\right).$
@@
The goal of this paper is to establish the message exchange protocol that supports the above learning goal.
Self-attention as communication protocol.
Suppose at time step $t$, each agent $i$ maintains a latent state $h_{t}^{i}\in\mathbb{R}^{n_{h}}$ (we use the horizontal vector notation) derived from its observations up to that point. To select an optimal action, an agent should incorporate information about other agents’ states. To facilitate such information sharing, we employ a self-attention mechanism Vaswani et al. (2017). Specifically, let $M^{k},M^{q},M^{v}\in\mathbb{R}^{n_{h}\times n_{h}}$ denote learnable projection matrices. Agent $i$ updates its representation by aggregating the states of all agents, where the contribution of agent $j$ is given by $\alpha^{i,j}_{t}h^{j}_{t}M^{v}$. The attention weight $\alpha^{i,j}_{t}$ is proportional to $\exp\left(h^{i}_{t}M^{q}(h^{j}_{t}M^{k})^{T}/\sqrt{n^{h}}\right)$, and is normalized such that $\sum_{j}\alpha^{i,j}_{t}=1$. This weighting reflects the relevance of agent $j$’s state to agent $i$ at time $t$.
The Transformer encoder introduced in Vaswani et al. (2017) augments self-attention with several additional components. These include (i) multi-head attention, which allows different subspaces of the state representations to be attended to in parallel; (ii) a feedforward MLP applied to the attention outputs; (iii) layer normalization applied after residual updates; and (iv) stacking multiple such blocks to form deep architectures. We adopt this Transformer encoder structure to enable effective and flexible information exchange among agents.
MACTAS.
@@
Intuitively, we understand the $f^{i}$ module as the “awareness” of the $i$-th agent. Based on what the agent is aware of, it determines its contribution to future discounted rewards, depending on the action it currently takes. In order to give a global view on the current state to the agent’s “awareness”, we introduce a communication mechanism. We divide the $f^{i}$ module after its recurrent part and feed its state $h^{i}_{t}\in\mathbb{R}^{n_{h}}$ to the module $g$, where the states of “awareness” of the agents are confronted. As a result, the $g$ module produces increments, $z^{i}_{t}\in\mathbb{R}^{n_{h}}$. The sum $z^{i}_{t}+h^{i}_{t}$ is then passed to the second MLP part of the $f^{i}$ module.
Consequently, the “awareness” of each agent is based not only on its observations and previous actions, but also on observations and previous actions of all other agents.
@@
  1. 1.
    It is fed with $n$ vectors and produces $n$ vectors, both of size $\dim(h^{i}_{t})$. The $i$-th input vector corresponds to the $i$-th output vector.
  2. 2.
@@
  3. 3.
    We use a separate Adam optimizer Kingma and Ba (2017) for the Transformer parameters; the rest of the architecture is trained with its originally proposed optimizer (typically RMSProp Tieleman and Hinton (2012)).
  4. 4.
    The weights of the output layer of the top Transformer module are initialized with zeros. Therefore, initially, this module does not affect the agents’ state of “awareness”.
A good performance of the Transformer in this role is expected. This architecture has been specifically optimized to transform a set of embeddings into another set of the same size, which is exactly what we require in the $g$ module.
Distributed deployment.
@@
A distributed implementation of a multi-agent system with our proposed module is also possible. Then, each agent needs to take on the part of processing that consists of refining $h^{i}_{t}$ and results in $z^{i}_{t}$. That requires exchanging messages with other agents as many times as there are Transformer layers in MACTAS, each time with $O(n)$ messages; $O(n^{2})$ for all agents. In training, the agents need to aggregate increments in their shared weights.